OpenStudy (anonymous):

Simplify the given equation. 17x - 6 + 3x - 5 = x + 11 + 4x

OpenStudy (anonymous):

Move all your terms with x over to the left, and your terms without x's to the right

OpenStudy (shamil98):

20x - 11 = 5x + 11 15x = 22 x = 22/15